OpenStudy (anonymous):

the length of a side of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length. if the perimeter of the triangle is 4 inches less than the perimeter of the rectangle, what are the dimensions of the rectangle?

OpenStudy (anonymous):

Solve the following equations for w and L: {w = L - 2, 3 L = 2 L + 2 w - 4} {w = 6, L = 8}